OpenStudy (anonymous):

Show all work in dividing 4t^2-1/15t+10 / 4t^2 - 4t + 1/ 3t^2+5t+2

OpenStudy (saifoo.khan):

can u please rewrite?

OpenStudy (anonymous):

4t^2-1/15t+10 (division bar) 4t^2-4t+1/3t^2+5t+2

OpenStudy (saifoo.khan):

are the figures correct?

OpenStudy (anonymous):

=[(2t+1)(2t-1)]/[5(3t+2)]/[(3t+2)(t+1)/(2t-1)(2t-1)] =[(2t+1)(t+1)]/(2t-1)

OpenStudy (mathteacher1729):

This thing is like the tower of Babel! \[\huge\frac{\frac{4t^2-1}{15t+10}}{\frac{4t^2-4t+1}{3t^2+5t+2}}\] STEP ONE Invert the denominator and multiply. \[ \huge \left( \frac{4t^2-1}{15t+10} \right) \cdot \left( \frac{3t^2+5t+2}{4t^2-4t+1}\right) \] STEP TWO: CRY. ;-p No, seriously, now just factor, individually, the numerator and denominator of each expression separately and cancel like a madman. Good luck!

OpenStudy (mathteacher1729):

Mertsj -- it won't be \(2(t+1)^2\) in the numerator because \((2t+1)(t+1) \neq 2(t+1)^2\)

OpenStudy (mertsj):